Q:- A string of length 'l' is fixed at both ends.It is vibrating in its 3rd overtone with maximum amplitude 'a'.The amplitude at a distance l/3 from one end is:


A. a


B. 0


C. (√3/2)a


D.a/2

Grade:12

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago
Dear Tapasranjan It is vibrating in its 3rd overtone then you will get nodes at x=0 ,l/4,2l/4,3l/4,l equation of motion is y= a sin(4?x/l)coswt put x=l/3 amplitude =a sin(4?l/3l) =a sin(4?/3) =a(v3/2) Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE & AIEEE preparation. All the best. Regards, Askiitians Experts Badiuddin
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

It is vibrating in its 3rd overtone
then you will get nodes at x=0 ,l/4,2l/4,3l/4,l
equation of motion is
y= a sin(4?x/l)coswt
put x=l/3
amplitude =a sin(4?l/3l)
=a sin(4?/3) =a(v3/2)

Thanks and Regards